Renee wants to test whether her number cube is fair. She decided that if the first rolled number keeps repeating, the cube may not be fair. She becomes suspicious of the cube landing on the same number consecutively if such an event happens to a fair number cube less than 1 time out of 50. Which is the shortest series of rolls that will make Renee think the cube is not fair?

A. NNNN

B. NNNNN

C. NNN...N } 49 times

D. NNN...N } 51 times
Answer #1

Here Renee wants to test whether her number if fair.She decided that if the first rolled number keeps repeating, the cube may not be fair.

She rolled the cube first time and get any number let say N.

Now, she rolled the cube second time and get again number N (i.e. NN), the probability of that happening = 1/6

Now, she rolled the cube third time and get again number N (i.e. NNN). The probability of that happening = 1/6 * 1/6 = 1/36

Now, again she rolled the cube fourth time and get again number N(i.e. NNNN). The probability of that happening = 1/6 * 1/6 * 1/6 = 1/216 , That is less than 1/50.

so here the shortest series of rolls that will make Renee think the cube is not fair = NNNN.

Option A is correct.